### v1

@@ Appendix C Calculus identities on manifolds
$=\left(\mathbf{P}\Delta_{\Gamma}\mathbf{u}-2(\mathbf{B}:\nabla_{M}\mathbf{u})\boldsymbol{\nu}-(\mathbf{u}\cdot\boldsymbol{\nu})\operatorname{div}_{\Gamma}\mathbf{B}-(\operatorname{div}_{\Gamma}\mathbf{B}\cdot\mathbf{u})\boldsymbol{\nu}-2\mathbf{B}\nabla^{T}_{M}(\mathbf{N}\mathbf{u})\,\boldsymbol{\nu}\right)_{i}.$
This proves (C.5). Identity (C.6) follows from multiplying (C.5) by $\mathbf{P}$ and recalling that $\mathbf{P}\boldsymbol{\nu}=\mathbf{0}$ and $\mathbf{P}\mathbf{B}=\mathbf{B}$. ∎
Proposition C.5 (symmetry of the covariant Hessian).
Assume $\Gamma$ is of class $C^{2}$. Then, for each $u\in C^{2}(\Gamma)$ the tensor field $\nabla_{\Gamma}\nabla_{\Gamma}u$ is symmetric.
Proof.
Recall definitions (2.1), (2.2) and (2.6), the facts that $\mathbf{P}\boldsymbol{\nu}=\mathbf{0}$ and $\mathbf{P}\mathbf{B}=\mathbf{B}$, and identity (2.7). Then, on one hand we have
@@
$(\nabla_{\Gamma}\nabla_{\Gamma}u)_{ij}$ $=(\mathbf{P}\nabla_{M}\nabla_{\Gamma}u)_{ij}=P_{ik}(\nabla_{M}\nabla_{\Gamma}u)_{kj}=P_{ik}\underline{D}_{j}\underline{D}_{k}u$
$=P_{ik}\underline{D}_{k}\underline{D}_{j}u+P_{ik}(\mathbf{B}\nabla_{M}u)_{k}\nu_{j}-P_{ik}(\mathbf{B}\nabla_{M}u)_{j}\nu_{k}=\underline{D}_{i}\underline{D}_{j}u+(\mathbf{B}\nabla_{M}u)_{i}\nu_{j}.$
On the other hand, again resorting to (2.7), we write
$(\nabla_{\Gamma}\nabla_{\Gamma}u)_{ji}=\underline{D}_{j}\underline{D}_{i}u+(\mathbf{B}\nabla_{M}u)_{j}\nu_{i}=\underline{D}_{i}\underline{D}_{j}u+(\mathbf{B}\nabla_{M}u)_{i}\nu_{j}-(\mathbf{B}\nabla_{M}u)_{j}\nu_{i}+(\mathbf{B}\nabla_{M}u)_{j}\nu_{i}=\underline{D}_{i}\underline{D}_{j}u+(\mathbf{B}\nabla_{M}u)_{i}\nu_{j},$
which shows the assertion. ∎

### v2

@@ Appendix C Riemann curvature tensor
$\langle R(\mathbf{a},\mathbf{b},\mathbf{v}),\mathbf{w}\rangle=-\int_{\Gamma}(\nabla_{\Gamma}\mathbf{v})\mathbf{b}\cdot\operatorname{div}_{\Gamma}(\mathbf{w}\otimes\mathbf{a})+\int_{\Gamma}(\nabla_{\Gamma}\mathbf{v})\mathbf{a}\cdot\operatorname{div}_{\Gamma}(\mathbf{w}\otimes\mathbf{b})-\int_{\Gamma}\Big[\nabla_{\Gamma}\mathbf{v}\,((\nabla_{M}\mathbf{b})\mathbf{a}-(\nabla_{M}\mathbf{a})\mathbf{b})\Big]\cdot\mathbf{w}.$ (C.3)
By means of the easily-verifiable identity $\operatorname{div}_{\Gamma}(\mathbf{a}\otimes\mathbf{b})=(\nabla_{M}\mathbf{a})\mathbf{b}+(\operatorname{div}_{\Gamma}\mathbf{b})\mathbf{a}$, we can equivalently rewrite (C.3) as
$\langle R(\mathbf{a},\mathbf{b},\mathbf{v}),\mathbf{w}\rangle$ $=\int_{\Gamma}(\nabla_{\Gamma}\mathbf{v})(\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a}):\nabla_{M}\mathbf{w}+\int_{\Gamma}\Big[\nabla_{\Gamma}\mathbf{v}\,\left((\nabla_{M}\mathbf{a})\mathbf{b}+(\operatorname{div}_{\Gamma}\mathbf{b})\mathbf{a}-(\nabla_{M}\mathbf{b})\mathbf{a}-(\operatorname{div}_{\Gamma}\mathbf{a})\mathbf{b}\right)\Big]\cdot\mathbf{w}$
$=\int_{\Gamma}(\nabla_{\Gamma}\mathbf{v})(\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a}):\nabla_{M}\mathbf{w}+\int_{\Gamma}\Big[\nabla_{\Gamma}\mathbf{v}\,\operatorname{div}_{\Gamma}(\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a})\Big]\cdot\mathbf{w}.$
Moreover, by using the identities $\nabla_{\Gamma}\mathbf{v}=\nabla_{M}\mathbf{v}+\boldsymbol{\nu}\mathbf{v}^{T}\mathbf{B}$ (cf. (2.7) and $\mathbf{B}\mathbf{w}=-(\nabla_{M}^{T}\mathbf{w})\boldsymbol{\nu}$ (cf. (2.6)), and the tangentiality of $\mathbf{w}$, we can further write
$\langle R(\mathbf{a},\mathbf{b},\mathbf{v}),\mathbf{w}\rangle=\int_{\Gamma}((\mathbf{v}^{T}\mathbf{B}\mathbf{b})\mathbf{B}\mathbf{a}-(\mathbf{v}^{T}\mathbf{B}\mathbf{a})\mathbf{B}\mathbf{b})\cdot\mathbf{w}\\ +\int_{\Gamma}\Big\{\left[\nabla_{M}\mathbf{v}(\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a})\right]:\nabla_{M}\mathbf{w}+\left[\nabla_{M}\mathbf{v}\,\operatorname{div}_{\Gamma}(\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a})\right]\cdot\mathbf{w}\Big\}.$
We now aim to prove that this last integral vanishes. For ease of notation, let $\mathbf{A}$ be the skew-symmetric tensor field given by $\mathbf{A}:=\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a}$. Adopting the Einstein summation convention, we have that
$(\nabla_{M}\mathbf{v})(\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a}):\nabla_{M}\mathbf{w}+\Big[\nabla_{M}\mathbf{v}\,\operatorname{div}_{\Gamma}(\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a})\Big]\cdot\mathbf{w}\\ =(\nabla_{M}\mathbf{v})\mathbf{A}:\nabla_{M}\mathbf{w}+\Big[\nabla_{M}\mathbf{v}\,\operatorname{div}_{\Gamma}\mathbf{A}\Big]\cdot\mathbf{w}=(\nabla_{M}\mathbf{v}\,\mathbf{A})_{ij}(\nabla_{M}\mathbf{w})_{ij}+(\nabla_{M}\mathbf{v}\,\operatorname{div}_{\Gamma}\mathbf{A})_{i}w_{i}\\ =(\nabla_{M}\mathbf{v})_{ik}A_{kj}(\nabla_{M}\mathbf{w})_{ij}+(\nabla_{M}\mathbf{v})_{ik}(\operatorname{div}_{\Gamma}\mathbf{A})_{k}w_{i}\\ =\underline{D}_{k}v_{i}A_{kj}\underline{D}_{j}w_{i}+\underline{D}_{k}v_{i}\underline{D}_{j}A_{kj}w_{i}=\underline{D}_{k}v_{i}\underline{D}_{j}(A_{kj}w_{i}).$ (C.4)
Now, let us notice that combining the integration-by-parts formula [11, Lemma B.5] with identity (2.3) and using a density argument, we easily arrive at the following identity:
@@
$\int_{\Gamma}\underline{D}_{i}u\,\underline{D}_{j}\varphi-\underline{D}_{j}u\,\underline{D}_{i}\varphi=\int_{\Gamma}\left\{(\mathbf{B}\nabla_{M}u)_{j}\nu_{i}-(\mathbf{B}\nabla_{M}u)_{i}\nu_{j}+\operatorname*{tr}(\mathbf{B})\left(\underline{D}_{i}u\,\nu_{j}-\underline{D}_{j}u\,\nu_{i}\right)\right\}\varphi,$ (C.5)
for each $u,\varphi\in C^{1}(\Gamma)$ and for each indices $i,j\in\{1,\dotsc,d+1\}$. Accordingly, applying this last identity to the right-hand side of (C.4), recalling the fact that $\mathbf{A}=\mathbf{a}\otimes\mathbf{b}-\mathbf{b}\otimes\mathbf{a}$ is skew-symmetric and that $\mathbf{a}$ and $\mathbf{b}$ are tangential, we arrive at
[…]
Hence, $\int_{\Gamma}\underline{D}_{k}v_{i}\underline{D}_{j}(A_{kj}w_{i})=0$ and so
$\langle R(\mathbf{a},\mathbf{b},\mathbf{v}),\mathbf{w}\rangle=\int_{\Gamma}((\mathbf{v}^{T}\mathbf{B}\mathbf{b})\mathbf{B}\mathbf{a}-(\mathbf{v}^{T}\mathbf{B}\mathbf{a})\mathbf{B}\mathbf{b})\cdot\mathbf{w},\qquad\mathbf{a},\mathbf{b},\mathbf{v},\mathbf{w}\in\mathbf{C}^{1}_{t}(\Gamma).$ (C.6)